The scores for a mathematics course seem to be normally distributed with mean 70 & standard deviation of 8. If the instructor wishes to give an A grade to 10% of the students, what should be the dividing line between an A grade and a B grade?


1
Expert's answer
2022-04-20T02:13:10-0400

Suppose that "X" is a random variable that corresponds to the score for a mathematics. Then, "P(X\geq a)=0.1". "a" corresponds to the dividing line between "A" and "B". We have: "P(X\geq a)=\int_{a}^{+\infty}f(x)dx", where "f(x)=\frac{1}{8\sqrt{2\pi}}e^{(-\frac12\left(\frac{x-70}{8}\right)^2)}". For "a=80.25" we get: "P(X\geq a)\approx0.100053". Thus, "80.25" is as a dividing line. (The answer is rounded up to two decimal places).